OpenStudy (anonymous):

An object is launched upward at 64 ft/sec from a platform that is 80 feet high. What is the object's maximum height if the equation of height (h) in terms of time (t) of the object is given by h(t) = -16t2 + 64t + 80? A. 272 B. 144 C. 1232 D. 616

OpenStudy (anonymous):

-16(2)^2+64(2)+80= -64+128+80=?

OpenStudy (anonymous):

144!

OpenStudy (anonymous):

Thanks much!

OpenStudy (anonymous):

Your welcome